# Nonparametric extensions of nuclear equations of state: probing the breakdown scale of relativistic mean-field theory

### v1

@@ IV Constraints on RMFT-breakdown with current astrophysical data
The symmetry energy and its slope at saturation are shown in Fig. 12 for the RMFT EoS set and hybridized priors transitioning at $p_{\rm t}/c^{2}=10^{13}\,\rm{g}/\rm{cm}^{3}$ and $10^{14}\,\rm{g}/\rm{cm}^{3}$. The posteriors are largely consistent with each other, though for higher transition pressures, and for the RMFT EoS set itself, the values of both $L$ and $J$ are better constrained. We find, for example, that $L=37.03^{+16.37}_{-14.42}\,$eV when we transition at $10^{13}\,\rm{g}/\rm{cm}^{3}$, and $L=47.87^{+11.15}_{-14.04}\,$eV when transitioning at $10^{14}\,\rm{g}/\rm{cm}^{3}$. The recovered value of $L$ is both larger and better constrained when we trust the RMFT up to higher densities, likely due to correlations between the large inferred maximum mass of neutron stars and the stiffness of matter near saturation under these models. Trusting the RMFT up to only $p_{\rm t}/c^{2}=10^{13}\,\rm{g}/\rm{cm}^{3}$, which corresponds to a density $\sim 1.5\,n_{0}$, we find no meaningful constraints on the symmetry parameters relative to the priors. This is an indication that under the model-agnostic prior, the properties of matter near saturation are not strongly correlated with the matter in the cores of astrophysical neutron stars Essick et al. (2021a, b).
The recovered values are consistent with existing results on the symmetry parameters, although they represent matter which is marginally softer near saturation than other studies. Reference Drischler et al. (2020a), using constraints from $\chi$EFT at N3LO, finds $J=31.7\pm 1.1\,\rm{MeV}$ and $L=59.8\pm 4.1\,\rm{MeV}$ at $1\sigma$ uncertainty. Additionally, per Fig. 2 of Ref. Drischler et al. (2020a) (based partly on Refs. Lattimer and Lim (2013); Lattimer and Steiner (2014)) this range is also found to be largely consistent with a broad range of theoretical, experimental, and observational constraints of the nuclear symmetry energy. Our results are also consistent with values implied by PREX-II Adhikari et al. (2021), though we do not consider that data here. For instance, Ref. Reed et al. (2021) used PREX-II results to infer $L=106\pm 61\,\rm{MeV}$ at $90\%$ credibility. Combining PREX-II and low-density $\chi$EFT information in the GP framework yields $L=53^{+14}_{-15}\,\rm{MeV}$ Essick et al. (2021a, b). Using a less flexible EoS model, though with a wider range of astrophysical and nuclear data including PREX-II, Ref. Biswas and Rosswog (2024) found $L=54^{+10}_{-10}\,\rm{MeV}$ at $90\%$.
@@ V Discussion
We have constructed hybrid models of the dense-matter EoS informed by an RMFT at low pressures and model-agnostic above a certain transition pressure. We compare multiple models with different transition pressures to astrophysical data, and probe the scale up to which a specific RMFT EoS is able to describe the dense-matter EoS. We first explore how inference behaves under simulated data. If the true EoS is consistent with the RMFT description, evidence will remain inconclusive, unable to rule out the RMFT at any scale. This behavior is expected from prior volume arguments. If the true EoS is inconsistent with any RMFT EoS, for example due to a phase transition, we recover strong evidence (Bayes factors $\gtrsim 10^{4}$) against models that are informed by the RMFT beyond the pressure of the phase transition. We further demonstrate that a breakdown can be identified with sufficiently many gravitational-wave observations in App. C, within expectations of upcoming observing campaigns.
Applying our models to real astrophysical observations, we find no evidence against RMFT EoSs up to arbitrarily high densities. Nonetheless, the inferred macroscopic neutron star properties differ depending on the pressure up to which the RMFT is trusted. This behavior elucidates how the breakdown pressure can be constrained with future data. For example, the inferred radius of a $1.4\leavevmode\nobreak\ M_{\odot}$ star depends on what pressure the RMFT is trusted up to. Conditioning up to a pressure of $1\times 10^{14}\,\rm{g}/\rm{cm}^{3}$, we find $12\lesssim R\lesssim 13\,\rm{km}$ at $\sim 90\%$ credibility, while conditioning up to only $3\times 10^{13}\leavevmode\nobreak\ \rm{g}/\rm{cm}^{3}$, yields $11-12.5\,\rm{km}$. This suggests that future constraints should be more informative about the breakdown scale if they reach $\sim 0.5\,\rm{km}$ precision and $R_{1.4}\sim 11.5\,\rm{km}$. Finally, as a note of caution, it is possible that biases emerge before evidence against a model appears. Therefore it is still important to consider other models of the EoS, including fully agnostic analyses. From this point of view, we see this work as complementary to existing studies which search for the breakdown of explicit models of low-density matter using astrophysical observations Biswas et al. (2021); Wouters et al. (2025).
@@ Acknowledgements.
We thank Thibeau Wouters for helpful comments on this manuscript. L.B. and A.H. thank Mark Alford for useful conversations. I.L. thanks Philippe Landry for insightful discussions. I.L. and K.C. acknowledge support from the Department of Energy under award number DE-SC0023101, and by a grant from the Simons Foundation (MP-SCMPS-00001470). L.B. and A.H. are partly supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award No. ##\##DE-FG02-05ER41375. A.H. furthermore acknowledges financial support by the UKRI under the Horizon Europe Guarantee project EP/Z000939/1. R.E. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC) through a Discovery Grant (RGPIN-2023-03346). The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Constructing the RMFT EoS distribution
In this Appendix, we discuss how we construct the distribution of RMFT EoSs. More details are available in Ref. Alford et al. (2022). We construct a particular RMFT EoS by making a choice for the RMFT nucleon-meson and meson-meson coupling constants; we choose these constants (which we will here refer to as parameters) so that the resulting EoS is consistent with the properties of isospin-symmetric nuclear matter and neutron matter near saturation. Since there is uncertainty in the properties of neutron matter, we sample these properties according to ab initio, $\chi$EFT, predictions. We then fit the RMFT parameters using a nonlinear, least-squares procedure, which allows us to produce a set of RMFT parameters which are consistent with the empirical saturation properties of symmetric nuclear matter, and ab initio constraints of neutron matter near saturation ($0.5-1.5\,n_{0}$). Therefore, we produce a distribution on the EoS where the only explicit uncertainty comes from uncertainty in the properties of neutron-rich matter near saturation.
@@
In contrast with isospin-symmetric matter, neutron matter is not self-bound and therefore cannot be probed in a lab, so we do not have analogous empirical constraints. Therefore, instead, we fit to ab initio $\chi$EFT calculations of neutron matter Tews et al. (2018). The $\chi$EFT calculation we use is based on local $\chi$EFT interactions at N2LO that were constructed in Refs. Gezerlis et al. (2013, 2014); Tews et al. (2016); Lynn et al. (2016). We choose this $\chi$EFT calculation because, compared to $\chi$EFT calculations at higher order like Refs. Drischler et al. (2019); Keller et al. (2021), Ref. Tews et al. (2018) employs a more conservative estimate of the error band. The $\chi$EFT results are used over the density range $0.5\,n_{0}$ to $1.5\,n_{0}$, i.e., above the crust-core transition and where $\chi$EFT is reliable Tews et al. (2018); Drischler et al. (2020b).
We sample the $\chi$EFT uncertainty band by creating a set of representative $\chi$EFT neutron matter EoSs using the Gandolfi-Carlson-Reddy (GCR) parametrization Gandolfi et al. (2012), which expresses the binding energy per nucleon $\mathcal{E}$ of neutron matter in the form
@@
We sample 200,000 EoSs generated from a wide range of GCR parameters $(a,b,\alpha,\beta)$, keeping the 1,109 EoSs that remain completely within the $\chi$EFT uncertainty band between $n_{B}=0.5\,n_{0}$ and $n_{B}=1.5\,n_{0}$. For each $\chi$EFT EoS, we fit the RMFT coupling constants, evaluating the $\chi$EFT EoS at 16 density points. A simultaneous fit to these points and the 12 density points of the phenomenological model for symmetric nuclear matter determines the RMFT coupling constants.
@@
Since the distribution of RMFT EoSs is influenced by the choice of initial guesses in the fitting procedure, the distribution we present here may not represent the full range of viable models consistent with low-energy nuclear physics for the chosen model Lagrangian. Nonetheless, the bulk of fit RMFT EoSs have maximum TOV masses below $2.0\,M_{\odot}$ regardless of the initial guess for the RMFT parameters. Therefore, we do not expect the bulk features of the RMFT EoS distribution to change based on the choice of initial guess, though the distribution of EoSs which exceed $2.0\,M_{\odot}$ is likely much more sensitive to this choice. Strategies to achieve particular distributions of RMFT EoSs are left to future work.
@@ Appendix B Efficiently Conditioning high-density GPs on Low-Density Theory
$p_{\mathrm{O}}(f_{a},f_{b})\sim\mathcal{N}\left([\mu_{a},\mu_{b}],\begin{bmatrix}C_{aa}&C_{ab}\\ C_{ba}&C_{bb}\end{bmatrix}\right)\,,$ (18)
@@ Appendix C Inferring a breakdown with gravitational-wave signals and pulsar timing measurements alone
We start with more details on the gravitational-wave injections. We use a binary population that is consistent with observations of merging neutron stars, sampling 50 sources from a representative population model which is uniform in component masses from 1.0 to $2.0\,M_{\odot}$. [^4] This is less than the TOV maximum mass for the EoSs we use for simulated data, assuming that the population of merging NSs is not limited by the TOV maximum mass; see e.g., Ref. Golomb et al. (2024) for a discussion. We assume this because high-mass neutron star mergers are essentially uninformative with respect to the EoS, and more importantly, high mass NSs are generally not distinguishable from black holes anyway in gravitational waves Farah et al. (2021); Fishbach et al. (2020); Essick and Landry (2020) We simulate signals in the Hanford and Livingston detectors Aasi et al. (2015) corresponding to the sampled sources into Gaussian noise at the level of A+ detector sensitivity a_p , placing the sources uniformly in comoving volume from 1 to 300 Mpc, which we expect to produce a substantial fraction of “informative” GW signals. We assume the neutron stars are spinning slowly, with dimensionless component spins isotropically distributed with magnitudes $|\vec{\chi}|<0.05$. We analyze all sources that are recovered with optimal signal-to-noise ratio (SNR) $>10.0$ using the bilby parameter estimation code Ashton et al. (2019) using the IMRPhenomPV2_NRTidalv2 waveform Dietrich et al. (2019). [^5] In general, selecting events based on the optimal SNR can induce biases Essick and Fishbach (2024). However, we do not change the mass, distance, or spin distribution, and the EoS contributes negligibly to selection effects. We then downsample the total set of analyzed sources in order to produce variable catalog sizes.

### v2

@@ IV Constraints on RMFT-breakdown with current astrophysical data
The symmetry energy and its slope at saturation are shown in Fig. 12 for the RMFT EoS set and hybridized priors transitioning at $p_{\rm t}/c^{2}=10^{13}\,\rm{g}/\rm{cm}^{3}$ and $10^{14}\,\rm{g}/\rm{cm}^{3}$. The posteriors are largely consistent with each other, though for higher transition pressures, and for the RMFT EoS set itself, the values of both $L$ and $J$ are better constrained. We find, for example, that $L=37.03^{+16.37}_{-14.42}\,$MeV when we transition at $10^{13}\,\rm{g}/\rm{cm}^{3}$, and $L=47.87^{+11.15}_{-14.04}\,$MeV when transitioning at $10^{14}\,\rm{g}/\rm{cm}^{3}$. The recovered value of $L$ is both larger and better constrained when we trust the RMFT up to higher densities, likely due to correlations between the large inferred maximum mass of neutron stars and the stiffness of matter near saturation under these models. Trusting the RMFT up to only $p_{\rm t}/c^{2}=10^{13}\,\rm{g}/\rm{cm}^{3}$, which corresponds to a density $\sim 1.5\,n_{0}$, we find no meaningful constraints on the symmetry parameters relative to the priors. This is an indication that under the model-agnostic prior, the properties of matter near saturation are not strongly correlated with the matter in the cores of astrophysical neutron stars Essick et al. (2021a, b).
[…]
@@ V Discussion
We have constructed hybrid models of the dense-matter EoS informed by an RMFT at low pressures and model-agnostic above a certain transition pressure. We compare multiple models with different transition pressures to astrophysical data, and probe the scale up to which a specific RMFT EoS is able to describe the dense-matter EoS. We first explore how inference behaves under simulated data. If the true EoS is consistent with the RMFT description, evidence will remain inconclusive, unable to rule out the RMFT at any scale. This behavior is expected from prior volume arguments. If the true EoS is inconsistent with any RMFT EoS, for example due to a phase transition, we recover strong evidence (Bayes factors $\gtrsim 10^{4}$) against models that are informed by the RMFT beyond the pressure of the phase transition. However, the strength of the evidence is sensitive to the parameters of the phase transition; in general, stronger transitions lead to stronger evidence of RMFT-breakdown with fewer observations. We further demonstrate that a breakdown can be identified with sufficiently many gravitational-wave observations in App. C, within expectations of upcoming observing campaigns.
Applying our models to real astrophysical observations, we find no evidence against RMFT EoSs up to arbitrarily high densities. When we assume no breakdown of the RMFT model, our inferred EoS distribution is qualitatively consistent with existing constraints on RMFT EoSs using astrophysical data Traversi et al. (2020); Malik et al. (2022); Huang et al. (2024); Zhou et al. (2023); Xia et al. (2024). Nonetheless, the inferred macroscopic neutron star properties differ depending on the pressure up to which the RMFT is trusted. This behavior elucidates how the breakdown pressure can be constrained with future data. For example, the inferred radius of a $1.4\penalty 10000\ M_{\odot}$ star depends on what pressure the RMFT is trusted up to. Conditioning up to a pressure of $1\times 10^{14}\,\rm{g}/\rm{cm}^{3}$, we find $12\lesssim R\lesssim 13\,\rm{km}$ at $\sim 90\%$ credibility, while conditioning up to only $3\times 10^{13}\penalty 10000\ \rm{g}/\rm{cm}^{3}$, yields $11-12.5\,\rm{km}$. This suggests that future constraints should be more informative about the breakdown scale if they reach $\sim 0.5\,\rm{km}$ precision and $R_{1.4}\sim 11.5\,\rm{km}$. Finally, as a note of caution, it is possible that biases emerge before evidence against a model appears. Therefore it is still important to consider other models of the EoS, including fully agnostic analyses. From this point of view, we see this work as complementary to existing studies which search for the breakdown of explicit models of low-density matter using astrophysical observations Biswas et al. (2021); Wouters et al. (2025).
@@ Acknowledgements.
We thank Thibeau Wouters for helpful comments on this manuscript. L.B. and A.H. thank Mark Alford for useful conversations. I.L. thanks Philippe Landry for insightful discussions. I.L. and K.C. acknowledge support from the Department of Energy under award number DE-SC0023101, and by a grant from the Simons Foundation (MP-SCMPS-00001470). L.B. and A.H. are partly supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award No. #\#DE-FG02-05ER41375. A.H. furthermore acknowledges financial support by the UKRI under the Horizon Europe Guarantee project EP/Z000939/1. R.E. is supported by the Natural Sciences & Engineering Research Council of Canada (NSERC) through a Discovery Grant (RGPIN-2023-03346). The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Constructing the RMFT EoS distribution
[…]
@@
[…]
@@
[…]
@@
Since the distribution of RMFT EoSs is influenced by the choice of initial guesses in the fitting procedure, the distribution we present here may not represent the full range of viable models consistent with low-energy nuclear physics for the chosen model Lagrangian. Nonetheless, the bulk of fit RMFT EoSs have maximum TOV masses below $2.0\,M_{\odot}$ regardless of the initial guess for the RMFT parameters. Therefore, we do not expect the bulk features of the RMFT EoS distribution to change based on the choice of initial guess, though the distribution of EoSs which exceed $2.0\,M_{\odot}$ is likely much more sensitive to this choice. Strategies to achieve particular distributions of RMFT EoSs, including direct sampling of RMFT parameters are left to future work.
@@ Appendix B Efficiently Conditioning high-density GPs on Low-Density Theory
We begin by discussing the particular choice of RMFT-informed and model-agnostic covariance kernels which are used in the construction. As stated in the main text, both the RMFT-informed and the model-agnostic EoS distributions in fact represent mixture models of Gaussian processes. The covariance matrices of the RMFT-informed distribution are induced directly by the distribution of RMFT EoSs and therefore have effectively zero parameters (being determined solely by how EoSs are assigned to subgroups to be fit by GPs). The agnostic distribution uses a wide array of possible kernel hyperparameters conditioned on a variety of EoSs, see Landry and Essick (2019); Essick et al. (2020a) for details. Different choices made in kernel construction may modify the details of the conclusions drawn here. However, we expect broad consistency because (1) the RMFT GP-EoS distribution is seen to effectively emulate the RMFT EoS distribution, and (2) the model agnostic EoS distribution has been found to have very broad support Landry and Essick (2019); Essick et al. (2020a, b, 2023) which would not change under small changes to any of the particular details of the construction.
@@
[…]
@@ Appendix C Inferring a breakdown with gravitational-wave signals and pulsar timing measurements alone
We start with more details on the gravitational-wave injections. We use a binary population that is consistent with observations of merging neutron stars, sampling 50 sources from a representative population model which is uniform in component masses from 1.0 to $2.0\,M_{\odot}$. [^4] This is less than the TOV maximum mass for the EoSs we use for simulated data, assuming that the population of merging NSs is not limited by the TOV maximum mass; see e.g., Ref. Golomb et al. (2024) for a discussion. We assume this because high-mass neutron star mergers are essentially uninformative with respect to the EoS, and more importantly, high mass NSs are generally not distinguishable from black holes anyway in gravitational waves Farah et al. (2021); Fishbach et al. (2020); Essick and Landry (2020) We simulate signals in the Hanford and Livingston detectors Aasi et al. (2015) corresponding to the sampled sources into Gaussian noise at the level of A+ detector sensitivity a_p , placing the sources uniformly in comoving volume from 1 to 300 Mpc, which we expect to produce a substantial fraction of “informative” GW signals. We assume the neutron stars are spinning slowly, with dimensionless component spins isotropically distributed with magnitudes $|\vec{\chi}|<0.05$. We analyze all sources that are recovered with optimal signal-to-noise ratio (SNR) >10.0>10.0 using the bilby parameter estimation code Ashton et al. (2019) using the IMRPhenomPV2_NRTidalv2 waveform Dietrich et al. (2019). [^5] In general, selecting events based on the optimal SNR can induce biases Essick and Fishbach (2024). However, we do not change the mass, distance, or spin distribution, and the EoS contributes negligibly to selection effects. We then downsample the total set of analyzed sources in order to produce variable catalog sizes.
@@
Appendix D Additional analysis with a weaker phase transition
We repeat our simulation analysis with the same RMFT EoS at low densities, but with a transition to a $c_{s}^{2}=0.7$ EoS at $2n_{0}$, with a $\Delta\varepsilon/\varepsilon$ of 0.3. The resulting EoS much more closely masquerades as an RMFT EoS. We show the analogous plot to Fig. 7 in Fig. 15. We find that such a change to the simulation EoS phase transition parameters substantially reduces the evidence against the RMFT model, however, we find the general trend seen in Fig. 7 still holds. Further, we show in Fig. 16 that when using the RMFT EoS set, the incorrect symmetry parameters are identified at the 90% level for this simulation EoS. Therefore, while astrophysical observations alone may struggle to distinguish such an observation, the additional inclusion of nuclear data from experiments and ab initio calculations should nonetheless allow the confident exclusion of the RMFT holding up to arbitrarily high densities.
Figure 15: Same as Fig. 7, but for a different choice of phase transition parameters.
Figure 16: Same as Fig. 8, but for a different choice of phase transition parameters.